A STUDY OF LINEWIDTH AND LINESHAPE: THE APPLICATION OF CONDON MODULATION TO PURE VIBRATIONAL TRANSITIONS OF SOLID HYDROGEN

\underline{MAN-CHOR CHAN}, , Institute for Astrophysics and Planetary Sciences, Ibaraki Universioty, 2-1-1 Bunkyo, Mito 310, Japan; YU ZHANG, AND TAKESHI OKA, Department of Chemistry, Department of Astronomy and Astrophysics, and the Enrico Fermi Institute, The University of Chicago, Chicago, IL 60637.

As first pointed out by Condon,^{a} a strong electric field can induce new features in the infrared spectra of molecules, which will follow the Raman selection rules. Following our observation of the pure vibrational Q_{1}(0) transition induced by the electric field of ionic species in solid parahydrogen, a new modulation technique with much improved sensitivity based on the Condon effect has been established.^{b} Using this technique, we have systematically studied the linewidth, frquency shift, and lineshape of the pure vibrational transitions of samples of parahydrogen crystals (with less than 0.07% of orthohydrogen) at various temperature. Results from this work provide information on the relaxation process in this system, which will lead to a better understanding of the origin of the sharpness of the transitions in solid hydrogens. In this paper, the analysis of our results will be presented.

aE. U. Condon, Phys. Rev. 41, 759 (1932).

bK. E. Kerr, T. Momose, W. P. Weliky, C. G. Gabrys, and T. Oka, Phys. Rev. Lett. 72, 3957 (1994).